# Multi-Objective Latent Space Optimization of Generative Molecular Design Models

### v2

@@ Introduction
Like other data-driven models, the initial capability of the generative molecular models to suggest novel molecules will be determined–at least to a certain extent–by the training data. In fact, the generated molecules are likely to reside in a similar chemical space as the molecules in the original training set, which may make it challenging to design new candidate molecules whose target attributes significantly exceed those of the original molecules–regardless of whether we perform molecular optimization in the latent space [5, 8, 9] or adopt a fine-tuning strategy [10, 11, 12, 13]. To mitigate this issue, a number of recent efforts aimed to improve the generative models by incorporating additional training data, generated either from experiments or simulations, where the goal is to ensure that these models are primed to suggest novel molecules with enhanced target properties that go beyond the initial molecules. Yang et al. [15] proposed an iterative retraining approach to improve the quality of the molecules sampled from the latent space of the generative model. In this approach, they pre-train the generative model jointly with a property predictor. The trained model is used to generate novel molecules from which a small batch of molecules with the best properties are selected using the predictor and added to the training dataset. The extended dataset is subsequently used for retraining to update the latent space to make it more amenable to produce better molecules with improved properties. In Liu et al. [16], the generator in a generative adversarial network (GAN) is iteratively updated under a chance-constrained optimization framework. They employ a validity function to guide the optimization of the property value within the region of grammatically correct input sequences or legitimate structures. A recent work by Tripp et al. [17] proposed a weighted retraining approach to reshape the latent space of a VAE to make it more sampling-efficient for producing novel molecules with improved properties. For this purpose, the weights assigned to the data points in the training dataset are determined by the rank of the corresponding molecules according to the objective function that evaluates the property of interest. Another recent iterative approach [18] applies a genetic algorithm (GA) along with domain knowledge to generate a new set of potentially improved candidates from the initial training data. The property values of these candidates are validated by high-throughput experiments or simulations. A deep neural network (DNN) property predictor is then retrained using a larger dataset augmented by the candidates generated by a GA. In contrast to the generative approaches [15, 16, 17], the candidates produced by a GA generally remain closer to the training data. As a result, the predictions from the DNN remain relatively reliable but the novelty of the produced candidates tends to be limited. Additionally, the (computational) cost of the experiments or simulations to assess the property of the novel candidates can be a significant burden when the candidate pool gets larger.
@@
In drug discovery applications, potential drug candidates always have to simultaneously meet multiple design criteria. For example, in addition to its capability to intervene with specific biomolecular target or mechanism, potential drug molecules are assessed based on various physio-chemical properties that contribute to their bioavailability. In the generative molecule design schemes discussed above, multi-objective optimization is typically handled via scalarization, which turns the problem into a simpler single-objective optimization problem. However, the weights for combining the multiple objective functions are often selected by the designer in an ad hoc manner, despite their importance in guiding the multi-objective optimization process. This may potentially lead to sub-optimal results, as the optimization process may be unintentionally dominated by a few objective functions, resulting in emphasizing certain attributes while ignoring others. Various other approaches have been also proposed to address multiple design criteria, where a notable example is RaionaleRL recently proposed by Jin et al. [19]. In this work, a policy gradient is applied to fine-tune a VAE for generating molecules from incomplete subgraphs that correspond to multi-property rationales. RationaleRL utilizes a property predictor to extract the subgraph rationale for each of the multiple target properties. From the set of extracted single property rationales, RationaleRL identifies the combined subgraph that meets the multiple property constraints and then uses the fine-tuned generative model to complete the molecular graph for the combined subgraph. However, a practical limitation of this approach is that the number of training samples that satisfy all properties of interest may be scarce, which exacerbates as the number of target design criteria increases. MArkov moleculaR Sampling (MARS) [20] proposed to address the multi-objective drug discovery problem by formulating the molecular design process as an iterative graph editing process. For this purpose, MARS defines the target distribution by combining the scoring functions of multiple properties and adopts Markov chain Monte Carlo (MCMC) sampling to identify high-scoring molecular candidates. However, as the target distribution is obtained by taking either the sum or the product of the multiple scoring functions, it faces similar shortcomings as other scalarization-based approaches discussed earlier.
@@
In this paper, we propose a novel multi-objective latent space optimization (MO-LSO) scheme that can effectively address the aforementioned limitations of existing generative models for molecular design. We extend the weighted retraining framework recently proposed in Tripp et al. [17] to equip it with the inherent capability to enhance the efficiency of sampling novel molecules in the latent space that simultaneously improve multiple target properties. This is achieved by ranking the molecules based on their Pareto optimality through non-dominated sorting (NDS), where the rankings are used both for generating improved molecules based on multiple design criteria to augment the training data and for determining the weights of the molecules in the (augmented) training set based on the relative importance. Our proposed MO-LSO scheme can naturally balance the trade-offs among multiple properties without any ad hoc scalarization that may potentially bias the optimization results. Furthermore, as the molecules are assessed based on their Pareto efficiency, our MO-LSO scheme scales very well as the number of design criteria increases and it is naturally equipped with the capability to handle the optimization of properties that may be highly correlated (or even redundant) or conflict one another. We show that our proposed MO-LSO scheme can effectively shift the latent space representation of the molecules based on multiple design criteria, thereby substantially enhancing the sampling efficiency of the generative model for suggesting novel molecules that simultaneously improve multiple properties. Furthermore, by applying it to the design of DRD2 inhibitors, we also demonstrate that the generative model optimized by the proposed MO-LSO scheme is able to produce highly promising molecules that outperform molecules designed by traditional QSAR prediction.
@@ Results
[…]
@@
As the baseline, we used the pretrained JT-VAE shared by Tripp et al. [17] and applied the proposed multi-objective weighted retraining scheme. Initially, we used the complete ZINC dataset[25] for weighted training of the baseline model, where the dataset was split into training ($218969$ molecules) and validation sets ($24333$ molecules) as in [17]. After each weighted retraining step, $250$ new molecules were randomly sampled, out of which the top $r=50$ candidates were selected based on the Pareto Front rank. The selected top candidates were used in the subsequent retraining stages. For details, please see the Methods section.
@@
Table S1 in the Supplementary Information shows the average property values for the molecules in the training data, molecules generated using the initial model (based on $1,000$ randomly generated molecules), and molecules generated by the retrained model (again based on $1,000$ randomly molecules) for different values of $k$. Additionally, we evaluated the structural diversity of the molecules generated by different versions of JT-VAE models, in order to assess the ability of a given model to learn, represent, and sample from a wider chemical space.
@@
Figure 3: Evolution of the Pareto front via latent space optimization of the generative model. The latent space of the JT-VAE has been jointly optimized to maximize logP and minimize SAS of the molecules suggested by the generative model. The scatter plots show the (logP, SAS) distribution of the molecules in the initial training dataset (column 1), molecules sampled in the latent space of the baseline model (column 2), and molecules suggested by the optimized model after iteration-1 (column 3), iteration-5 (column 4), iteration-10 (column 5). The plots in the top row show the trends for the case when the complete training dataset was used, while the bottom row shows the trend when a reduced dataset was used. The results show that the Pareto front gradually shifts towards the desired direction (i.e., bottom right for larger logP and smaller SAS).
Next, we investigated the ability of the weighted retraining scheme to propose high-scoring molecules when such molecules are absent in the training dataset. For this purpose, we first removed the top $20\%$ molecules – selected based on their Pareto front rank – from the training data that is used to train the baseline model [17]. Since the original baseline model has seen the complete dataset during its training, for each property pair, we trained a separate baseline model based on the reduced dataset that does not contain the top 20% molecules with the highest Pareto front rank for the given property pair. In all cases, a learning rate of $0.0007$ was used with a batch size of $32$ for $30$ epochs for model training. The scatter plots in Fig. 3 show the progression of the two properties logP and SAS, where the latent space of the JT-VAE is optimized for the given property pair with $k=10^{-5}$. The top row of Fig. 3 shows the evolution of the Pareto front when the complete dataset was used. On the other hand, the bottom row in Fig. 3 shows the trends for the reduced dataset which does not contain the top 20% molecules. As can be seen in the bottom row of Fig. 3, although the initial dataset does not contain many molecules with logP greater than $4$ and SAS lower than $2$, the iterative multi-objective weighted retraining still manages to effectively push the latent space towards a desirable region that contains high-scoring molecules, where the trends are similar to the case when the complete dataset is used.
@@
Figure 4 shows the results for the multi-objective weighted retraining with $k=10^{-6}$ for the property pair DRD2 and SAS. The leftmost scatter plot shows the distribution of the training data after removing the top 20% molecules. The horizontal axis shows the probability of inhibition (the higher the better) and the vertical axis shows the SAS (a lower score corresponds to better synthesizability). The second subplot (from the left) shows the property distribution of $1,000$ molecules randomly sampled in the latent space of the baseline model. Next, the third, fourth, and fifth (i.e., rightmost) subplots depict the property distribution of the molecules after the $1^{st}$, $5^{th}$ and $10^{th}$ weighted retraining cycle, respectively. While the initial training data are absent of DRD2 inhibitory molecules, we can see that a relatively large number of inhibitory molecules after the $10^{th}$ weighted retraining. Moreover, the SAS distribution of the sampled molecules becomes more skewed towards smaller values as expected. We have repeated similar experiments for two other property pairs (DRD2, logP) and (DRD2, NP score), which all showed similar trends. Details of these simulation results can be found in Table S2 of Supplementary Information.
Figure 4: Transition of DRD2 and SAS towards optimum direction while starting with no DRD2 active training samples. The first two scatter plots represent the training data and the molecules generated from the baseline model respectively. In rest of the jointplots, the molecules from learned model after $1^{st}$, $5^{th}$ and $10^{th}$ weighted retraining for $k=10^{-6}$ are shown in the objective space.
@@
Figure 6: Properties of a top molecule predicted by the optimized generative model. The properties of a top molecule (compound A) suggested by the JT-VAE, whose latent space was optimized by the proposed method. Six parameters – POLAR (Polarity), INSOLU (Insolubility), INSATU (Instauration), FLEX (Rotable bond flexibility), LIPO (Lipophilicity), and SIZE (Molecular Weight) – are shown. We can see that the suggested compound is within the colored zone, which corresponds to the physio-chemical space suitable for oral bioavailability.
While in silico results do not often translate in activity in vivo, they are often used to select what molecules get made and their priority for testing, so the results of this effort and the computational chemistry evaluation produced “high priority compounds” that are very different from the native ligand and the molecules on the market. This ability to in silico explore vast amounts of novel chemistry space while still generating compounds with desirable properties is likely to lead to the types of cost and time reductions in discovery that have been promised by the AI/ML community. The actual compounds predicted have desirable properties (Figure 6 shows the properties for compound A as predicted by SwissADME [26]). Additionally, the binding of compound A (Figure 7 (a)) shows that the typical aromatic interaction between W386 and the ligand in this case the $6$ member ring on the ligand as well as the potential for salt bridges from the ligand N’s to the O’s on D114. Of more interest is the two cyclopropane groups, which allow the compounds to have hydrophobic interactions with the largely hydrophobic mouth of the pocket, which includes residues 389, 392, 184, 189, 416, 412, etc. (Figure 7 (b)) without having undesirable steric effects that bulkier ring systems may present. The potential desolvation effects from these residues while still allowing for the protein to have a closed confirmation have the potential to allow the compounds to be selective and active at low concentrations. As previously mentioned, this molecular series would not have been ranked high by computational or medicinal chemists based on how they looked, but once the poses were reviewed and the ease of the synthetic routes evaluated it is generally agreed that the series is a high priority.
@@
Figure 8: Effect of Bayesian optimization on the progression of the model in the objective space: The first and second rows represent the scatter plots for random selection and Bayesian optimization strategy respectively. In each row, the objective space are for the training data, the molecules from the baseline model and retrained model after $1^{st}$, $5^{th}$ and $10^{th}$ weighted retraining for $k=10^{-3}$.
In previous sections, we generated novel molecules through random sampling in the latent space of the trained generative model, among which the top molecules were selected to augment the training data for the next cycle of weighted retraining. Although we adopted random sampling in order to demonstrate how the proposed multi-objective weighted retraining can enhance the sampling efficiency of the generative model for multiple target molecular properties, the overall efficacy of the GMD can be further improved by leveraging more sophisticated optimization techniques – such as Bayesian optimization (BO) or genetic algorithms (GA) – to optimize the molecules in the latent space. To demonstrate this, we utilized BO to optimize molecules in the latent space of the generative model for logP and SAS. A single objective function was defined through scalarization, where logP (to be maximized) was penalized by SAS (to be minimized) after standardizing both values based on the mean and standard deviation of the respective values in the initial dataset. Based on the expected improvement acquisition function, we generate $50$ molecules and only unique samples are added to the training dataset.

### v3

@@ Introduction
Like other data-driven models, the initial capability of the generative molecular models to suggest novel molecules will be determined–at least to a certain extent–by the training data. In fact, the generated molecules are likely to reside in a similar chemical space as the molecules in the original training set, which may make it challenging to design new candidate molecules whose target attributes significantly exceed those of the original molecules–regardless of whether we perform molecular optimization in the latent space [5, 8, 9] or adopt a fine-tuning strategy [10, 11, 12, 13]. To mitigate this issue, a number of recent efforts aimed to improve the generative models by incorporating additional training data, generated either from experiments or simulations, where the goal is to ensure that these models are primed to suggest novel molecules with enhanced target properties that go beyond the initial molecules. Yang et al. [15] proposed an iterative retraining approach to improve the quality of the molecules sampled from the latent space of the generative model. In this approach, they pre-train the generative model jointly with a property predictor. The trained model is used to generate novel molecules from which a small batch of molecules with the best properties are selected using the predictor and added to the training dataset. The extended dataset is subsequently used for retraining to update the latent space to make it more amenable to produce better molecules with improved properties.
A similar effort from Iovanac et al. [16] utilizes the grammar variational autoencoder [17], jointly trained with a linear predictor network to sample new molecules along the latent dimension corresponding to the targeted property region. The new molecules go through further screening to be used with the training dataset to retrain the generative model iteratively. In Liu et al. [18], the generator in a generative adversarial network (GAN) is iteratively updated under a chance-constrained optimization framework. They employ a validity function to guide the optimization of the property value within the region of grammatically correct input sequences or legitimate structures. A recent work by Tripp et al. [19] proposed a weighted retraining approach to reshape the latent space of a VAE to make it more sampling-efficient for producing novel molecules with improved properties. For this purpose, the weights assigned to the data points in the training dataset are determined by the rank of the corresponding molecules according to the objective function that evaluates the property of interest. Another recent iterative approach [20] applies a genetic algorithm (GA) along with domain knowledge to generate a new set of potentially improved candidates from the initial training data. The property values of these candidates are validated by high-throughput experiments or simulations. A deep neural network (DNN) property predictor is then retrained using a larger dataset augmented by the candidates generated by a GA. In contrast to the generative approaches [15, 18, 19, 16], the candidates produced by a GA generally remain closer to the training data. As a result, the predictions from the DNN remain relatively reliable but the novelty of the produced candidates tends to be limited. Additionally, the (computational) cost of the experiments or simulations to assess the property of the novel candidates can be a significant burden when the candidate pool gets larger.
@@
In drug discovery applications, potential drug candidates always have to simultaneously meet multiple design criteria. For example, in addition to its capability to intervene with specific biomolecular target or mechanism, potential drug molecules are assessed based on various physio-chemical properties that contribute to their bioavailability. In the generative molecule design schemes discussed above, multi-objective optimization is typically handled via scalarization, which turns the problem into a simpler single-objective optimization problem. However, the weights for combining the multiple objective functions are often selected by the designer in an ad hoc manner, despite their importance in guiding the multi-objective optimization process. This may potentially lead to sub-optimal results, as the optimization process may be unintentionally dominated by a few objective functions, resulting in emphasizing certain attributes while ignoring others. Various other approaches have been also proposed to address multiple design criteria, where a notable example is RaionaleRL recently proposed by Jin et al. [21]. In this work, a policy gradient is applied to fine-tune a VAE for generating molecules from incomplete subgraphs that correspond to multi-property rationales. RationaleRL utilizes a property predictor to extract the subgraph rationale for each of the multiple target properties. From the set of extracted single property rationales, RationaleRL identifies the combined subgraph that meets the multiple property constraints and then uses the fine-tuned generative model to complete the molecular graph for the combined subgraph. However, a practical limitation of this approach is that the number of training samples that satisfy all properties of interest may be scarce, which exacerbates as the number of target design criteria increases. MArkov moleculaR Sampling (MARS) [22] proposed to address the multi-objective drug discovery problem by formulating the molecular design process as an iterative graph editing process. For this purpose, MARS defines the target distribution by combining the scoring functions of multiple properties and adopts Markov chain Monte Carlo (MCMC) sampling to identify high-scoring molecular candidates. However, as the target distribution is obtained by taking either the sum or the product of the multiple scoring functions, it faces similar shortcomings as other scalarization-based approaches discussed earlier. To generate molecules with multiple target properties, Feng et.al [23] utilized a Langevin diffusion process as a stochastic generator of latent embeddings within a pre-trained autoencoder model. This stochastic generator, primarily governed by reference molecules selected based on multiple properties of interest, can be seen as an alternative to other optimization strategies in latent space.
@@
In this paper, we propose a novel multi-objective latent space optimization (MO-LSO) scheme that can effectively address the aforementioned limitations of existing generative models for molecular design – specifically, their limited capability for extrapolation in a multi-objective fashion beyond the molecular property space seen during training. We extend the weighted retraining framework recently proposed in Tripp et al. [19] to equip it with the inherent capability to enhance the efficiency of sampling novel molecules in the latent space that simultaneously improve multiple target properties. This is achieved by ranking the molecules based on their Pareto optimality through non-dominated sorting (NDS), where the rankings are used both for generating improved molecules based on multiple design criteria to augment the training data and for determining the weights of the molecules in the (augmented) training set based on the relative importance. Our proposed MO-LSO scheme can naturally balance the trade-offs among multiple properties without any ad hoc scalarization that may potentially bias the optimization results. Furthermore, as the molecules are assessed based on their Pareto efficiency, our MO-LSO scheme scales very well computationally as the number of design criteria increases (empirically demonstrated up to three objectives in this study) and it is naturally equipped with the capability to handle the optimization of properties that may be highly correlated (or even redundant) or conflict one another. We show that our proposed MO-LSO scheme can effectively shift the latent space representation of the molecules based on multiple design criteria, thereby substantially enhancing the sampling efficiency of the generative model for suggesting novel molecules that simultaneously improve multiple properties. Furthermore, by applying it to the design of DRD2 inhibitors, we also demonstrate through in silico analysis that the generative model optimized by the proposed MO-LSO scheme is able to produce highly promising molecules that outperform known DRD2 inhibitory molecules.
@@ Results
In this study, we considered the pairwise optimization of the following molecular properties: (i) water-octanol partition coefficient (logP), (ii) synthetic accessibility score (SAS), (iii) natural product-likeness score (NP score) [25], and (iv) the probability of inhibition against the dopamine receptor D2 (DRD2) [26]. We aimed to maximize logP, NP score, and DRD2 inhibition property. On the other hand, we aimed to minimize SAS, as a lower SAS indicates better synthesizability of a given molecule. For the computation of logP and SAS, the RDKit package [27] was used. We adopted the method in [25] for assessing the NP score. The probability of inhibition against DRD2 was estimated by using an ML surrogate model, whose details are given in the Methods section.
@@
As the baseline model, we used the pretrained JT-VAE shared by Tripp et al. [19] and applied the proposed multi-objective weighted retraining scheme. Initially, we used the complete ZINC dataset[28] for weighted training of the baseline model, where the dataset was split into training ($218969$ molecules) and validation sets ($24333$ molecules) as in [19]. After each weighted retraining step, $250$ new molecules were randomly sampled, out of which the top $r=50$ candidates were selected based on the Pareto Front rank. The selected top candidates were used in the subsequent retraining stages. For details, please see the Methods section.
@@
Table S1 in the Supplementary Information shows the average property values for the molecules in the training data, molecules generated using the initial model (based on $1000$ randomly generated molecules), and molecules generated by the retrained model (again based on $1000$ randomly molecules) for different values of $k$. Additionally, we evaluated the structural diversity (Table S1) of the molecules generated by different versions of JT-VAE models, in order to assess the ability of a given model to learn, represent, and sample from a wider chemical space.
@@
To further demonstrate the effectiveness of our proposed approach MO-LSO, we performed additional experiments for simultaneous optimization of two or three molecular properties and compared the results against the scalarization baseline of [19] and a Markov molecular sampling scheme for multi-objective drug discovery, called MARS [22]. Specifically, the latter approach involves training a molecular generative model from scratch targeting multiple properties of interest whereas our approach starts with a generative model that is trained in a self-supervised fashion without the molecular property values of the training samples. However, through iterative retraining via our approach, the generative model succeeds in outperforming MARS for quite a few property combinations (Table S6). Furthermore, we investigated the impact of retraining on the JT-VAE’s molecule reconstruction performance. The results obtained from these experiments are summarized and discussed in Supplementary Information.
@@
Figure 3: Evolution of the Pareto front via latent space optimization of the generative model. The latent space of the JT-VAE has been jointly optimized to maximize logP and minimize SAS of the molecules suggested by the generative model. The scatter plots show the (logP, SAS) distribution of the molecules in the initial training dataset (column 1), molecules sampled in the latent space of the baseline model (column 2), and molecules suggested by the optimized model after iteration-1 (column 3), iteration-5 (column 4), iteration-10 (column 5). The plots in the top row show the trends for the case when the complete training dataset was used, while the bottom row shows the trend when a reduced dataset was used. The results show that the Pareto front gradually shifts towards the desired direction (i.e., bottom right for larger logP and smaller SAS) resulting in a larger hypervolume (HV) of the Pareto front dominated property space in both cases.
Next, we investigated the ability of the weighted retraining scheme to propose high-scoring molecules when such molecules are absent in the training dataset. For this purpose, we first removed the top $20\%$ molecules – selected based on their Pareto front rank – from the training data that is used to train the baseline model [19]. Since the original baseline model has seen the complete dataset during its training, for each property pair, we trained a separate baseline model based on the reduced dataset that does not contain the top 20% molecules with the highest Pareto front rank for the given property pair. In all cases, a learning rate of $0.0007$ was used with a batch size of $32$ for $30$ epochs for model training. The scatter plots in Fig. 3 show the progression of the two properties logP and SAS, where the latent space of the JT-VAE is optimized for the given property pair with $k=10^{-5}$.
In each plot, we also show the hypervolume of the property space that is dominated by the Pareto front. The hypervolume is computed with respect to the average property values of the molecules in the complete training dataset. The top row of Fig. 3 shows the evolution of the Pareto front when the complete dataset was used. On the other hand, the bottom row in Fig. 3 shows the trends for the reduced dataset which does not contain the top 20% molecules. As can be seen in the bottom row of Fig. 3, although the initial dataset does not contain many molecules with logP greater than $4$ and SAS lower than $2$, the iterative multi-objective weighted retraining still manages to effectively push the latent space towards a desirable region that contains high-scoring molecules, where the trends are similar to the case when the complete dataset is used. This is also illustrated by the larger hypervolume achieved by the optimized model compared to the initial pre-trained model in both cases.
@@
Figure 4 shows the results for the multi-objective weighted retraining with $k=10^{-6}$ for the property pair DRD2 and SAS. The leftmost scatter plot shows the distribution of the training data after removing the top 20% molecules. The horizontal axis shows the probability of inhibition (the higher the better) and the vertical axis shows the SAS (a lower score corresponds to better synthesizability). The second subplot (from the left) shows the property distribution of $1,000$ molecules randomly sampled in the latent space of the baseline model. Next, the third, fourth, and fifth (i.e. rightmost) subplots depict the property distribution of the molecules after the $1^{st}$, $5^{th}$ and $10^{th}$ weighted retraining cycle, respectively. While the initial training data are absent of DRD2 inhibitory molecules, we can see that a relatively large number of inhibitory molecules after the $10^{th}$ weighted retraining. Moreover, the SAS distribution of the sampled molecules becomes more skewed towards smaller values as expected. We have repeated similar experiments for two other property pairs (DRD2, logP) and (DRD2, NP score), which all showed similar trends. Details of these simulation results can be found in Table S2 of Supplementary Information.
Figure 4: Transition of DRD2 and SAS towards optimum direction while starting with no DRD2 active training samples. The first two scatter plots represent the training data and the molecules generated from the baseline model respectively. In rest of the jointplots, the molecules from learned model after $1^{st}$, $5^{th}$ and $10^{th}$ weighted retraining for $k=10^{-6}$ are shown in the objective space. The hypervolume (HV) in each plot indicates the volume of the property space that is dominated by the Pareto front. As we iterate the weighted retraining, the resulting hypervolume tends to increase.
@@
Figure 6: Properties of a top molecule predicted by the optimized generative model. The properties of a top molecule (compound A) suggested by the JT-VAE, whose latent space was optimized by the proposed method. Six parameters – POLAR (Polarity), INSOLU (Insolubility), INSATU (Instauration), FLEX (Rotable bond flexibility), LIPO (Lipophilicity), and SIZE (Molecular Weight) – are shown. We can see that the suggested compound is within the colored zone, which corresponds to the physio-chemical space suitable for oral bioavailability [29].
While in silico results do not often translate in activity in vivo, they are often used to select what molecules get made and their priority for testing, so the results of this effort and the computational chemistry evaluation produced “high priority compounds” that are very different from the native ligand and the molecules on the market. This ability to in silico explore vast amounts of novel chemistry space while still generating compounds with desirable properties is likely to lead to the types of cost and time reductions in discovery that have been promised by the AI/ML community. The actual compounds predicted have desirable properties (Figure 6 shows the properties for compound A as predicted by SwissADME [29]). Additionally, the binding of compound A (Figure 7 (a)) shows that the typical aromatic interaction between W386 and the ligand in this case the $6$ member ring on the ligand as well as the potential for salt bridges from the ligand N’s to the O’s on D114. Of more interest is the two cyclopropane groups, which allow the compounds to have hydrophobic interactions with the largely hydrophobic mouth of the pocket, which includes residues 389, 392, 184, 189, 416, 412, etc. (Figure 7 (b)) without having undesirable steric effects that bulkier ring systems may present. The potential desolvation effects from these residues while still allowing for the protein to have a closed confirmation have the potential to allow the compounds to be selective and active at low concentrations. As previously mentioned, this molecular series would not have been ranked high by computational or medicinal chemists based on how they looked, but once the poses were reviewed and the ease of the synthetic routes evaluated it is generally agreed that the series is a high priority. For illustrative purpose, we performed the ADMET analysis for the top candidate for DRD2. In practice, if one desires to select the optimal candidate leads for the properties of interest, a more comprehensive screening like the one performed in [23] may be required.
@@
Figure 8: Effect of Bayesian optimization on the progression of the model in the objective space: The first and second rows represent the scatter plots for random selection and Bayesian optimization strategy respectively. In each row, the objective space are for the training data, the molecules from the baseline model and retrained model after $1^{st}$, $5^{th}$ and $10^{th}$ weighted retraining for $k=10^{-3}$. Higher increase in hypervolume (HV) for Bayesian optimization strategy shows its effectiveness over random selection.
In previous sections, we generated novel molecules through random sampling in the latent space of the trained generative model, among which the top molecules were selected to augment the training data for the next cycle of weighted retraining. Although we adopted random sampling in order to demonstrate how the proposed multi-objective weighted retraining can enhance the sampling efficiency of the generative model for multiple target molecular properties, the overall efficacy of the GMD can be further improved by leveraging more sophisticated optimization techniques – such as Bayesian optimization (BO), genetic algorithms (GA), and particle swarm optimization (PSO) [6] – to optimize the molecules in the latent space. To demonstrate this, we utilized BO to optimize molecules in the latent space of the generative model for logP and SAS. A single objective function was defined through scalarization, where logP (to be maximized) was penalized by SAS (to be minimized) after standardizing both values based on the mean and standard deviation of the respective values in the initial dataset. Based on the expected improvement acquisition function, we generate $50$ molecules and only unique samples are added to the training dataset. Compared to the random generation strategy, the Bayesian optimization approach is computationally expensive since it requires training the GP surrogate model as well as optimizing the acquisition function. However, the BO approach is more sample efficient as it requires fewer property predictions to find the best data points to augment the training dataset. On the other hand, the random generation approach requires a larger number of evaluations (250 in our case) to select the best $50$ candidates for data augmentation.